### v2

# A kiloparsec-scale ordered magnetic field in a galaxy at z=5.6

A kiloparsec-scale ordered magnetic field in a galaxy at z=5.6
Magnetic fields are prevalent on almost all astronomical scales, but their importance in different systems and over cosmic time is yet to be understood. Our current knowledge of the evolution of magnetic fields is limited by scarce observations in the distant Universe, where galaxies have recently been found to be more evolved than most model predictions. In this study, we conducted rest-frame 131 m full-polarisation observations of dust emission in a strongly lensed dusty star-forming galaxy, SPT034652, at , when the Universe was only 1 Gyr old. Dust grains can become aligned with local magnetic fields, resulting in the emission of linearly polarised thermal infrared radiation. Our observations have revealed a median polarisation level of per cent with a variation of per cent across the regions with polarisation detection, similar to that of local starburst galaxies. The polarised dust emission is patchy and mostly overlaps with the [C ii] emission at a velocity of about 150 km s-1, and extends over 3 kiloparsecs. The bimodal distribution of field orientations, their spatial distribution, and the connection with the cold gas kinematics further emphasise the complexity of the magnetic environment in this galaxy and the potential role of mergers in shaping its magnetic fields. Such early formation of ordered galactic magnetic fields also suggests that both small-scale and large-scale dynamos could be efficient in early galaxies. Continued observations of magnetic fields in early galaxies, as well as expanding surveys to a wider galaxy population, are essential for a comprehensive understanding of the prevalence and impact of magnetic fields in the evolving Universe.
Key Words.:
galaxies: high-redshift – galaxies: distances and redshifts – galaxies: formation – galaxies: starburst – submillimeter: galaxies1 Introduction
Magnetic fields are ubiquitous in the multi-phase interstellar medium (ISM), and circum-galactic medium (CGM) of galaxies (Beck & Wielebinski 2013; Planck Collaboration et al. 2015; Han 2017; Borlaff et al. 2023), but their origin remains a mystery. The development of magnetic fields is generally thought to start from seed fields and then be amplified through various dynamos (see the recent reviews by Subramanian 2019; Brandenburg & Ntormousi 2023). The seed fields could be primordial, either a relic of the early Universe (during the inflation and phase transition) or generated along with the earliest cosmic structure formation, with a strength G (Durrer & Neronov 2013; Subramanian 2019). Beyond this, along with the formation of proto-galaxies, Population-III stars, supernovae, and early active galactic nuclei (AGN) could also deposit up to G seed fields into the ISM (e.g. Hanayama et al. 2005; Xu et al. 2008; Beck et al. 2013; Attia et al. 2021). These seed fields can then be amplified through the small-scale dynamo (also called fluctuation dynamo, turbulent dynamo, Brandenburg & Subramanian 2005), which converts the turbulent kinetic energy of the gas into magnetic field energy until the medium at a given scale saturates (i.e. reach equipartition). In turn, the amplified turbulent fields from the small-scale dynamo undergo ordering processes, such as - dynamo (Ruzmaikin et al. 1988; Beck 2015; Brandenburg & Ntormousi 2023), to form large-scale magnetic fields spanning sub-kiloparsec (kpc) to tens of kpc scales. Theoretically, the amplification and ordering timescale are sensitive to the model assumptions, which make the observation of early magnetic fields crucial to discriminating different models (e.g. Arshakian et al. 2009; Rodrigues et al. 2019).
The current most explored method is Faraday rotation; it measures the rotation of the polarisation angle when a background linearly polarised light passes through a magneto-ionic medium. By statistically measuring the Faraday rotation in a sample of high-redshift polarised quasars with and without intervening galaxies, G regular field (parallel to the line-of-sight) have been reported in these intervening galaxies up to redshifts of about 2 (e.g. Bernet et al. 2008, 2013; Kronberg et al. 2008; Farnes et al. 2014, 2017). However, the total rotation measures integrate all the contributions along the line of sight, including the ISM/CGM of the quasars, intervening material, and Galactic magnetic fields. Different samples or different models of the Galactic magnetic fields can sometimes lead to controversial results (Oren & Wolfe 1995; Kim et al. 2016; Lan & Prochaska 2020; Malik et al. 2020). Recently, Mao et al. (2017) circumvented this difficulty by measuring the differential Faraday rotation between two sub-images of a lensed quasar, whose light passes through different parts of a foreground late-type galaxy at . Assuming the differential rotation comes mainly from the different field strengths within the foreground lensing galaxy, the Faraday depth indicates G-level regular magnetic fields around the foreground galaxy. With improved sensitivity of the radio telescopes, the Faraday rotation measurements have been applied to an increasingly larger number of intervening galaxies towards the sightlines of the quasars with different impact parameters (e.g. Böckmann et al. 2023). Meanwhile, in addition to distant quasars, the fast radio bursts (FRBs) have become another promising background polarised source (e.g. Mannings et al. 2023). All these progresses show a growing community of mapping magnetic fields in galaxies on various spatial scales and at different cosmic epochs. More importantly, the next-generation radio telescopes, e.g. the Square Kilometre Array (SKA) and the next-generation Very Large Array (ngVLA)(Beck et al. 2019; Heald et al. 2020), can further test these results and push the measurements to an earlier cosmic time.
Polarised dust is another promising magnetic field tracer across cosmic time. Dust grains are not perfectly spherical; when they are exposed to anisotropic radiation fields, the radiation absorbed by the surface of the dust will create net radiative torque (RAT) to rotate the grains along the minor axis. Dust grains are also paramagnetic, which causes the rotating dust to generate an internal magnetic moment that ends up aligned with the orientation of the local magnetic field. The final configuration is that the dust grain’s minor axis aligns with the local magnetic fields by Larmor precession to minimise the magnetic moment of inertia (Lazarian & Hoang 2007; Hoang & Lazarian 2008, 2016). This mechanism is known as B-RAT, which is the most accepted mechanism driving the grain alignment in the ISM (Andersson et al. 2015; Lopez-Rodriguez & Tram 2024). In nearby galaxies, far-infrared (FIR) polarimetric observations have been the most efficient way to probe the geometry of the projected magnetic fields in the cold phase of the interstellar medium (Lopez-Rodriguez et al. 2021, 2022a, 2022b; Pattle et al. 2021; Lopez-Rodriguez 2023). To differentiate the magnetic fields measured with different techniques, hereafter, we will use “ordered” fields to represent the inferred magnetic fields from polarised dust emission, which has 180 degrees of ambiguity, and use “regular” for fields with a well-defined direction.
Dust is ubiquitous among star-forming galaxies (SFGs). The advent of submillimetre/millimetre (submm/mm) telescopes has pushed the detections of dust emission back to the epoch of cosmic reionisation, benefiting from the negative- correction (e.g. Laporte et al. 2017; Tamura et al. 2019; Inami et al. 2022). Meanwhile, hundreds of the most bright dusty SFGs have been confirmed to be strongly lensed systems, with their apparent flux densities and spatial scales being amplified by a foreground galaxy or galaxy cluster (e.g. Reuter et al. 2020; Urquhart et al. 2022). The lens amplification can facilitate the detection of faint polarised dust emission and boost the spatial resolution. In addition, polarised thermal dust emission is not affected by Faraday rotation. All these advantages make dust polarisation a promising option to map the galactic magnetic fields across cosmic time with modern submm/mm interferometers. The current highest-redshift detection of polarised dust emission in a lensed galaxy, 9io9, at has been recently reported by Geach et al. (2023). The dust polarisation fraction in 9io9 is of the order of 1 per cent and the polarised dust emission extends over 5 kpc along the galaxy’s disk. The polarisation fraction in 9io9 is comparable to nearby star-forming galaxies, which is intriguing considering that 9io9 has an order of magnitude higher gas mass and star formation rate (SFR). The ordered magnetic fields are generally aligned with the cold gas disk, favouring a rapid formation of ordered magnetic fields. Such a success also opens a new window to explore magnetic fields in distant galaxies.
In this paper, we conduct full-polarisation Atacama Large Millimeter/Submillimeter Array (ALMA) dust continuum observation of 2 at an even higher redshift, z=5.56 ( Gyr after the Big Bang, the Cosmic Dawn), to map its galactic magnetic fields with polarised dust emission. SPT034652 is the brightest and most intensely star-forming galaxy discovered in the 2,500 deg2 survey of the South Pole Telescope (Vieira et al. 2010; Hezaveh et al. 2013). Lens modelling shows that SPT034656 has been magnified by a foreground galaxy (at ) with an amplification factor of and a half-light radius of kpc (Spilker et al. 2016). Correcting for the lens amplification, SPT034652 is an intrinsic hyper-luminous infrared galaxy (HyLIRG, L) and has a SFR around (Ma et al. 2016). Given its compact size, SPT034652 has one of the most extreme SFR surface densities (Ma et al. 2016). In addition, SPT034652 has a highly enriched ISM, dominated by warm and dense molecular gas (Apostolovski et al. 2019; Litke et al. 2022). The total flux density of SPT034652 at m is mJy (Reuter et al. 2020), which makes it the best target to detect and resolve the polarised dust emissions back to the Cosmic Dawn.
The paper is organised as follows. In section 2, we describe the basic information about the full-polarisation observations and data reduction. In section 3, we derive the linear polarisation properties of the dust emission and the geometry of the inferred magnetic fields. In section 4, we compare the observed results with the nearby galaxies and discuss possible physical origins of the ordered magnetic fields, their formation timescale, and their implications for early galaxy evolution. Finally, we summarise our work in section LABEL:sec:conclusion. A flat CDM cosmology with , , and km s-1 Mpc-1 is assumed throughout this work.
2 Observation and data reduction
| Date | # of Ant. | Resolution | PWVa | Noise Levelc | |
| (arcsec) | (mm.) | (min.) | (Jy/beam) | ||
| 2022-08-24 | |||||
| 2022-08-30 |
The full-polarisation observations of SPT034652 were conducted in ALMA Band 7 under ALMA project ID: 2021.1.00458.S, (P.I. Jianhang Chen). The data were obtained on the 24th and 30th of August 2022, with 283(116) min and 192(77) min observation(on-source) time, respectively. The details about the observing conditions are listed in Tab. 1. In both observations, the coverage of parallactic angles is wider than 120 degrees, which delivers independent full-polarisation calibration.
The data were calibrated with Common Astronomy Software Applications (CASA) package (ver.6.2.1, The CASA Team et al. 2022). We first followed the standard calibration pipeline of the two parallel polarisations, XX and YY, wherein J05194546 and J02535441 were used to calibrate the bandpass and phase variation, respectively. The flux of all the targets was calibrated from the flux calibrator J05194546, with an uncertainty of about 5 per cent. Then, the calibrated data were split for polarisation calibration. We followed the official template to apply the polarisation calibration. For both sessions, the blazar J04230120 was used as the polarisation calibrator to correct the instrumental polarisation leakage. The best-fit polarisation fraction of J04230120 is around 4.70.1 per cent, consistent with the long-term monitoring of its polarisation status of 3.2-5.1 per cent at the same period 222https://www.alma.cl/ skameno/AMAPOLA/J0423-0120.flux.html. The measured uncertainty in polarisation angle is 0.2 deg. The instrumental uncertainty of the linear polarised fraction is about 0.1 per cent and 1 deg in terms of polarisation angle from ALMA technical handbook (Cortes et al. 2023).
After the polarisation calibration, we used task tclean within CASA to make the images and cubes. We first created image cubes of the Stokes I (the total intensity) to identify possible spectral lines. The cubes were created with the ‘hogbom’ deconvolver and natural weighting. We did not identify any strong lines () within our spectral windows. Then, we created the full Stokes images with the ‘clarkstokes’ deconvolver and natural weighting to maximise the signal-to-noise ratio (S/N). All the clean processes stop when the residuals reach the root mean squared (RMS) level of 12 Jy beam-1 with a beam size of ( kpc at =5.6). The RMS achieved for the polarised intensity is 8.5 Jy beam-1.
We also used the archive ALMA Band 7 observation of [C ii] from project ID: 2013.1.01231 (PI: Dan Marrone) to quantify the kinematic states of SPT034652. The detailed analyses of [C ii] have been reported in Litke et al. (2019). We reduced the data following the ”scriptforPI.py” with CASA (ver.4.2.2). Before imaging, the continuum was subtracted from the visibility using uvcontsub with the ”fitorder” = 1. Different from the ”robust” weighting scheme used in Litke et al. (2019), we adopted a natural weighting scheme when cleaning the images with tclean shipped with CASA (ver.6.2.1). The natural weighting delivers better sensitivity and better matches the resolution of our full-polarisation dust observations. We also binned the channel width of the [C ii] datacube to 100 km s-1.
3 Results and analysis
In Fig. 1, we show the Stokes , and images of SPT034652. The observation probes the dust polarisation at the rest-frame wavelength of 131 m. The polarised intensity is calculated by
Following the Nyquist theorem, we re-sampled the Stokes maps with an aperture that is half the size of the clean beam. Within each sampling, we calculated the median values of and . The median value is more robust than the mean value in SPT034652 due to the general low signal-to-noise ratio (S/N) of each Stokes parameter. Changing to the mean values will not change our results. Then, we derived the polarisation fraction within each sampling using:
where , , are the median value of the corresponding Stokes parameter within each sampling. Before calculating the position angle of the magnetic fields , we masked the regions with . Then, the position angle is calculated from:
the additional is added because the projected magnetic field orientation is perpendicular to the measured polarisation angle of dust emission.
The inferred magnetic fields with Nyquist sampling are shown in Figs. 1 and 2, overlaid on the total and polarised intensity map, respectively. The histograms of the polarisation fraction, , and magnetic field position angle, , are also shown in Fig. 2. The vertical dashed and dotted lines indicate the median and mean of polarisation fractions among all the samplings. The median and mean values of polarisation fractions both show per cent, with 0.4 per cent variations across the individual polarisation measurements. The measured polarisation fraction is close to the measurement in 9io9, where a mean fraction of per cent is reported at rest-frame 350 m and extends over 5 kpc parallel to the galactic disk (Geach et al. 2023). The position angles of the magnetic fields in SPT034652 have a large variation, ranging from degrees, and show a bimodal distribution. The vertical dash-dotted lines show the two peaks of the bimodal distribution at 28 degrees and 78 degrees, with median polarisation fractions of and per cent, respectively.
To recover the dust emission in the source plane, we used a pixellated lensing reconstruction technique developed by Warren & Dye (2003) and used extensively in the field with the public available software (PyAutoLens, Nightingale & Dye 2015; Nightingale et al. 2021; Dye et al. 2018, 2022). This method makes no assumption about the source structure, allowing complex source features to be recovered. To better guide the lens modelling, we masked the central area with 0.05 arcsec and the outer area with 0.18 arcsec. The reconstructed source models are shown in Fig. 3 and the best-fit lens parameters are summarised in Tab. 2. The Stokes of SPT034652 is reconstructed to a single elongated and relatively compact region of emission in the source plane, which is consistent with the existing lens models (e.g. Spilker et al. 2015; Litke et al. 2019; Jones et al. 2019a). The recovered spatial structure of the polarised emission is offset roughly 300-600 pc from the peak of Stokes and only present in the north and east sides of the lensed galaxy. The position angles of the field line change gradually from the north side to the east side, which favours the presence of ordered magnetic fields in these regions. However, due to the faintness of the polarised dust emission, we can only marginally recover its structure in the source plane by modelling the region with possible signals. We are exploring different ways to model lensed full-polarisation data, including the modelling in the visibility space. The details will be reported in the forthcoming paper by Liu et al. (in prep.). The physical resolution of our observation varies at different galactic positions. The region with the polarisation detection has the highest amplification, with a mean amplification factor of , which corresponds to pc in the source plane with our angular resolution. After correcting the lens amplification, the polarised dust extends over 3 kpc in the source plane.
| Parameters | Definition | Value |
| R.A. of mass centre | 03:46:41.43 | |
| R.A. of mass centre | -52:05:06.36 | |
| Redshift of lensing galaxy | 0.9 | |
| Redshift of source galaxy | 5.565 | |
| Elliptical component x | ||
| Elliptical component y | ||
| Einstein radius (arcsec) | ||
| Slope | ||
| Elliptical components of shear in x | ||
| Elliptical components of shear in x |
4 Discussion
In section 3, we have reported the properties of polarised dust in SPT034652, which indicates that kpc-scale ordered magnetic fields are already present at the Cosmic Dawn. Here, we focus on several key questions that can be addressed from these observations: (1) How different are the magnetic fields between the distant and local galaxies? (2) What are the main drivers in ordering galactic magnetic fields? (3) What is the formation timescale of magnetic fields? (4) How could the early large-scale magnetic fields affect galaxy evolution?
4.1 Polarisation fraction and morphology of magnetic fields
Owing to the wavelength coverage and the spatial resolution provided by ALMA, we can observe similar rest-frame FIR and reach comparable spatial resolutions as the dust polarisation observations in the nearby Universe, especially compared with what High-resolution Airborne Wideband Camera Plus (HAWC+) has done in the SALSA Legacy Program (Lopez-Rodriguez et al. 2022b). In Fig. 4, we compare the measured dust polarisation fraction in SPT0346-52 with the measurements from the SALSA Legacy Program. The rest-frame wavelength probed in SPT034652 is at 131 m, averaged over a physical resolution about 700 pc, comparable to the spatial resolution of SALSA Legacy Program at 154 m (250 pc-1.5 kpc, Lopez-Rodriguez et al. 2022b). A typical mean resolved polarisation fraction of per cent has been reported among all the galaxies in the SALSA survey, dropping to a lower fraction of per cent if considering only starburst galaxies (SFR). We also include the dust polarisation measurements of 9io9 (Geach et al. 2023). The similar polarisation fraction between SPT034652 and the local starbursts is surprising, even though there are two orders of magnitude differences in SFR and more than 12 Gyr difference in cosmic time.
In SPT034652, the polarised dust emission is more patchy. We only observe polarised dust emission in a part of the galaxy and, interestingly, it does not overlap with the peak of total dust intensity (see Fig. 2). In Fig. 5, we compare the spatial distribution of polarisation with the kinematics from [C ii]. The region with the polarisation detection is also the region with the lowest velocity gradient and velocity dispersion. In the channel maps of [C ii] emission, the polarisation is mostly associated with a single velocity component near . This result is different from the configuration of polarised dust emission in 9io9, where the polarised dust is largely overlapped with the total dust emission and the inferred fields line are aligned with cold molecular gas disk (Geach et al. 2023). Such a difference suggests a possible different ordering process (see more discussion in section 4.2).
4.2 Physical origin of the ordered magnetic field
Various processes could lead to the formation of large-scale ordered magnetic fields, including the dynamo in galactic disks (Beck et al. 1994; Beck 2007), the stretching and compressing of existing fields by galaxy mergers, and galactic inflows/outflows (Chyży & Beck 2004; Drzazga et al. 2011; Basu et al. 2017; Jones et al. 2019b; Lopez-Rodriguez et al. 2022a; Lopez-Rodriguez 2023; Ledos et al. 2024). In this section, we will discuss several possible mechanisms to explain the observed magnetic fields in SPT034652.
4.2.1 Disk or merger driven?
High-resolution spectral line observations of SPT034652 have resolved a large velocity gradient ( 600 km s-1) of the cold gas in the ISM (Spilker et al. 2015; Litke et al. 2019). Due to the compact size of SPT034652 in the source plane, it is still unclear whether it’s a rotating disk or an ongoing major merger. Litke et al. (2019) showed that a lensing reconstruction of the high-resolution [C ii] emission revealed two distinct components, both spatially and kinematically. However, based on the observation of water absorption, Jones et al. (2019a) concluded that a compact disk with a strong molecular outflow gives a better explanation. In this subsection, we discuss both scenarios in generating the ordered magnetic fields.
The formation of disks is the most discussed channel to develop galactic ordered magnetic fields (Beck et al. 1994; Pakmor et al. 2017, 2024). The presence of a galactic disk can provide continuous differential rotation to support the dynamo (Beck et al. 2019). Initially, dynamo was not favoured in the early Universe, as most of the early galaxies were thought to be irregular and turbulent. This picture may need to be revised, however, given the apparent prevalence of disks in the early Universe (Förster Schreiber et al. 2018; Rizzo et al. 2023; Kuhn et al. 2023, Lee et al. (in prep.)). If galaxies can develop disks efficiently during the first Gyr of the cosmic time, it is reasonable to expect large-scale ordered magnetic fields in the disks of early galaxies (Arshakian et al. 2009; Geach et al. 2023).
In SPT034652, the rich observations allow a quantitative diagnosis of dynamo associated with a galactic disk. Following the simplified model by Arshakian et al. (2009), we can define the dynamo number:
where the is the scale height of the disk, is the angular velocity of the disk; and is the maximum velocity and velocity dispersion of the disk; is the maximum radius of the disk. Within a galactic disk, the dynamo is effective only when the dynamo number is larger than the critical value , which is close to 7 in Milky Way-like disk galaxies. The from our lens model is about 1.5 kpc (see also Fig. 3). Assuming a typical disk scale height of 300 pc (Geach et al. 2023), the is in the order of . Considering the complex velocity field (merger+outflows), it is not straightforward to estimate the without some biased assumptions. In the most extreme case, assuming the velocity gradient is solely produced by a galactic disk, the maximum rotational velocity will be km s-1. We adopted the median velocity dispersion of 140 km s-1 from the region with the lowest velocity gradient, which is also the region with polarisation detection (see also Fig. 5). Then, the is about 2.5, which predicts , well below the (Ruzmaikin et al. 1988; Arshakian et al. 2009, see also the discussion in Geach et al. (2023)). Therefore, even if there is a disk in SPT034652, it will not be very effective in ordering the galactic magnetic fields. However, if the dynamic cold disk () could exist at such early cosmic time (e.g. Rizzo et al. 2020; Neeleman et al. 2020; Lelli et al. 2021; Rowland et al. 2024), the dynamo number could be large enough to allow efficient dynamo.
Galaxy mergers can compress and stretch magnetic fields, creating large-scale ordered magnetic fields. This process has been well demonstrated in the local major merger prototype – NGC4038/39 (the Antennae) (Chyży & Beck 2004; Basu et al. 2017). In the cores of the two merging galaxies of the Antennae system, the field lines are mostly dominated by turbulent fields. However, in the tidal tail, the magnetic fields are regular up to kpc, which is likely generated by stretching the initial disk fields in the pre-merger galaxies (Basu et al. 2017). A 9 kpc ordered magnetic field is also reported by means of magnetically aligned dust grains, which also supports a compressed and stretched magnetic field due to the merger activity (Lopez-Rodriguez et al. 2022a). The presence of ordered magnetic fields has also been confirmed in a sample of interacting galaxies observed with synchrotron polarisation (at 4.86 and 1.4 GHz) (Drzazga et al. 2011), which all support tidal interactions can effectively magnetise the mergers’ surroundings and produce large-scale ordered field structures up to tens of kpc in the ISM of the merged galaxies. Besides, the developed kpc-scale ordered magnetic fields have a strength several times larger than the magnetic fields in disks from non-interacting galaxies (Drzazga et al. 2011; Beck & Wielebinski 2013). Similar intense field amplification and creation of large-scale magnetic fields with galaxy mergers have also been noticed in magneto-hydrodynamic simulations, which is crucial for angular momentum transportation and subsequent morphological evolution of the merged galaxies (Whittingham et al. 2021, 2023).
In SPT034652, the polarised dust emission is offset from the total dust emission, similar to the polarised dust distribution in Antennae galaxies (Lopez-Rodriguez et al. 2022a), which is likely due to the disruption of ordered fields by the merger-trigger starbursts. In addition, the bimodal distribution of also favour the merger scenario. The position angles of the field lines depend only on the ratio of Stokes and , which have received identical amplification and averaging from the strong lensing. The presence of two groups of in the image plane indicates that the source plane should also have at least two magnetic field components, i.e., and in Fig. 2. For a merging system like the Antennae galaxies, several configurations of magnetic field can be formed, i.e. the existing fields on each galaxy, the fields connecting both galaxies, and the newly stretched lines in the tidal tail. Alternatively, a disk could also show a bimodal distribution of , but it needs to be highly inclined towards the sightline (such as NGC7331 in Borlaff et al. 2023), and it cannot explain why the polarised dust mostly overlapped with a single velocity component. Besides, the dynamo predicts regular magnetic fields with axisymmetric position angle in the disk, which is not fully consistent with the field structures from the lens reconstruction, where ordered fields have only been observed in the north and east sides of the lensed galaxy. The compact size of SPT034652 favours a late-stage merger. It supports the possibility that galactic magnetic fields have already been established in the two merging galaxies in SPT034652, and the merger-triggered starburst makes the B-field more turbulent and leaves mainly the ordered magnetic fields in the outskirts, such as the north and east sides of the lensed galaxy in the source plane. Therefore, the galaxy merger provides the most natural explanation for the formation of ordered magnetic fields in SPT034652.
4.2.2 Outflow driven?
A strong outflow is present in SPT034652 based on the high-resolution dust continuum and H2O() observations (Jones et al. 2019a). It has a velocity up to 500 km s-1 and a substantial outflow rate, 100-900 M⊙ yr-1, likely driven by the galactic wind powered by the compact starbursting core. The galactic wind can distribute amplified magnetic fields to a larger volume (Brandenburg et al. 1993; Arámburo-García et al. 2021), thus can also explain the ordered field in the outskirts of SPT034652.lactic disks (Lopez-Rodriguez et al. 2021).
However, the outflow can only partially explain the observations. Firstly, outflow can provide a distinct component along with the components from the host galaxy (Lopez-Rodriguez 2023), which can easily explain the bimodal distribution of in SPT034652, but the recovered field lines in the source plane seem to disfavour the bipolar configuration if SPT034652 is a disk galaxy. Besides, while the outflow scenario is compatible with the ordered field overlapped with a single velocity component, the relatively low-velocity dispersion and velocity gradient in the region with polarization detection also disfavours the outflow as the main driver. Future higher resolution follow-up of SPT034652 could provide better constraints.
4.2.3 AGN driven?
It is also worth noting the possible contribution from the AGN in SPT034652. Powerful jets from AGN play an important role in magnetising the intracluster medium (ICM, Dubois et al. 2009; Xu et al. 2009). Meanwhile, the large-scale shocks induced by the jet and the AGN-driven outflow could also produce large-scale shear to order the turbulent fields. In the less extreme radio-quiet AGNs, the ubiquitous presence of compact radio jets (e.g. Girdhar et al. 2022; Rivera et al. 2023) also indicates their possible contribution to the development of magnetic fields in the ISM. In SPT034652, however, the presence of AGN is still elusive. The spectrum energy distribution (SED) analysis suggests a negligible AGN contribution to the total energy output, which is supported by the non-detection of X-ray and faint radio emission (Ma et al. 2016). In nearby AGN host galaxies, only the radio-loud AGN show highly polarised FIR fraction (5–11 per cent) in the core region; the radio-quiet AGN are mostly unpolarised (Lopez-Rodriguez et al. 2023). Therefore, even though there is a possible embedded AGN in SPT034652, our results support that it plays a negligible role in ordering the galactic magnetic fields. Future surveys targeting the AGN host galaxies are needed to test AGN’s importance in magnetising the ISM.
4.2.4 Caveats
In addition to the physical origins, the spatial resolution could also affect the observed polarisation. For the FIR polarisation, the polarised dust emission is density weighted by each individual molecular cloud, whose size is well below the current observation. A coarser spatial resolution thus introduces stronger beam averaging, which could counterbalance the polarised signal if the field lines are highly tangled or change orientation rapidly within each beam (Girichidis 2021). Gravitational lensing further complicates the situation by giving a different physical resolution along and perpendicular to the critical curve. Nevertheless, it represents the best way to mitigate the beam averaging. A robust polarisation signal has been observed in 9io9 (Geach et al. 2023), where the observation also benefits from a large lens amplification () to probe the dust polarisation down to 600 pc. In SPT034652, the regions with the polarisation are also those with the highest amplification, which indicates the non-detection in the other regions could be attributed to a combination of more turbulent magnetic fields and stronger beam averaging. This is also likely to be the reason for the non-detection in two other unlensed submm galaxies observed by Kade et al. (2023), where the physical resolution is worse than kpc.
In summary, a merger event provides the most likely explanation of the ordered magnetic fields in SPT034652, but we cannot fully rule out the possible contribution from outflows and a compact galactic disk. It is also possible that all these processes may work simultaneously to amplify and order the magnetic fields. Therefore, future high-resolution follow-ups will be the key to distinguishing different scenarios and alleviating the beam averaging.
4.3 Formation timescales of galactic magnetic fields
Theoretically, the amplification via small-scale dynamos is efficient, capable of amplifying the seed magnetic fields to G turbulent magnetic within a timescale of 100 Myr (Schleicher et al. 2010; Pakmor et al. 2014; Rieder & Teyssier 2016; Martin-Alvarez et al. 2022). However, the large-scale dynamos may need timescales of several Gyr to create kpc-scale ordered fields (e.g. Wang & Abel 2009; Kotarba et al. 2009; Arshakian et al. 2009; Rodrigues et al. 2019).
The starburst nature of SPT034652 allows efficient small-scale dynamos in the ISM. In local star-forming galaxies, the total magnetic field strength correlates with the SFR surface density, i.e. with to be around 0.3 (e.g. Heesen et al. 2014), indicating a close bond between the magnetic fields and star formation activities. In SPT034652, the SFR surface density is about . Such concentrated SFR normally indicates a highly turbulent ISM (Elmegreen & Scalo 2004; Bournaud et al. 2010), which may lead to the efficient amplification of turbulent magnetic fields. Meanwhile, the presence of turbulent magnetic fields has been used to connect the two SFR tracers: the total FIR luminosity and non-thermal radio luminosity (i.e. FIR-radio correlation) (Condon 1992; Yun et al. 2001; Bell 2003). The FIR-radio correlation has been observed up to by stacking analysis and, interestingly, it shows only mild variations with the redshift (e.g. Ivison et al. 2010; Magnelli et al. 2015; Delvecchio et al. 2021; Yoon 2024). This close correlation suggests that amplified magnetic fields may already be present in the star-forming galaxies up to , which supports the claim that small-scale dynamo efficiently amplifies the turbulent magnetic fields in the early star-forming galaxies. Fig. 4 shows the magnetic field strength as a function of redshift and total SFR. Due to the selection effects, the current two measurements of dust polarisation at , SPT0346-52 and 9io9, are both extreme starburst galaxies, which biased our results to the most extreme SFGs in the early Universe.
As we have mentioned before, the excess of Faraday rotation in the quasar absorber and fast radio burst support the presence of regular magnetic fields in galaxies up to (e.g. Bernet et al. 2008, 2013; Kronberg et al. 2008; Mannings et al. 2023). The detection of dust polarisation in 9io9 shows the ordered magnetic fields are present at z=2.6 (Geach et al. 2023). The redshift of SPT0346-52 is 5.6, thus the detection of the dust polarisation indicates the development of kpc-scale ordered galactic magnetic fields can be less than 1 Gyr, which puts a strong constraint on the ordering timescale. The current results do not allow us to directly test the dynamo. However, if the stretching and compressing during a merger is the main reason for the observed ordered magnetic fields (see the discussion in section 4.2), it indicates the two merging galaxies may have already developed ordered fields before the final merger, which indicates the ordering process is already ongoing during the first Gyr of the cosmic history. Therefore, our results suggest that the ordering process from large-scale dynamos could be efficient in early galaxies. Similar to the Faraday rotation, a more robust sample covering different galaxy types, redshifts, and environments is required to confirm this claim.
The detection of magnetic fields at Cosmic Dawn can also constrain the amplification of their strength. However, estimating the strength from dust polarisation is non-trivial. The most widely used method, the DavisChandrasekharFermi (DCF) method (Davis & Greenstein 1951; Chandrasekhar & Fermi 1953), has mostly been used in the Galactic molecular clouds and star-forming regions (Crutcher 2012; Pattle et al. 2022), where the spatial resolution is able to capture the turbulence of the small-scale magnetic fields. On the galactic scale, modified DCF has been applied to the outflow region of M82 and the central molecular zone of the Milky Way (Lopez-Rodriguez et al. 2021; Guerra et al. 2023), but all of them require the spatial resolution comparable to the turbulence scale. The physical resolution of our observation at the location with polarised dust emission has a physical resolution of pc in the source plane. However, the turbulent scale in starburst galaxies injected by supernovae is found to be about 50-500 pc (Elmegreen & Scalo 2004), which is smaller than our observation and thus prevents us from estimating the strength of magnetic fields reliably.
As we detected dust emission polarisation, it indicates that there is an ordered magnetic field associated with the cold phase of the ISM. The maximum magnetic field strength that this cold phase could reach is the equipartition energy level with the turbulent kinetic energy of the cold phase. Therefore, we can derive an upper limit by assuming an equipartition between the turbulent kinetic and magnetic energies (see also Geach et al. 2023). The equipartition field strength is formulated as
where is the volume density of the ISM, and is the velocity dispersion of the gas. In SPT0346-52, the observed [C ii] line dispersion at the region with polarisation detection is about 140 km s-1 (see also Fig. 5). We assume a preferred gas density of 50 cm-3 from the photodissociation regions (PDR) modelling (Litke et al. 2022). We estimate the equipartition turbulent magnetic field strength to be G. We emphasise that this is only a theoretical upper limit for the turbulent magnetic fields, with the assumption that the turbulence dominates the line width. If the line width is dominated by regular rotation or the velocity difference of the possible two merging galaxies, the equipartition field strength can be several times lower, which will bring the field strength closer to the local starburst galaxies, either measured with dust polarisation (Lopez-Rodriguez et al. 2021; Lopez-Rodriguez 2023) or radio synchrotron observations (Lacki & Beck 2013). It should be noted that the magnetic field strength in the cold phase of the interstellar medium is different from that of the warm and diffuse medium measured at radio wavelengths (Martin-Alvarez et al. 2023). The field strength in the cold and dense medium could be larger than that from the warm and diffuse medium, as the field strength correlates with the gas volume density. Nonetheless, our calculation supports that highly amplified magnetic fields could be developed in very young star-forming galaxies.
4.4 Implications for early galaxy evolution
Magnetic fields have been widely accepted as an essential part of galaxy evolution, but their importance is still under debate, especially on galactic scales. For instance, to affect the gas collapse/infall on a free-fall timescale, the ordered field strength needs to be G (e.g. Ceverino et al. 2012), which is 10100 larger than most observations. However, Krumholz & Federrath (2019) summarised the effects of magnetic fields in various star-formation processes, concluding that they are not the dominant player in controlling star formation inside galaxies, but they could play a larger indirect role by the interaction with stellar feedback and cosmic rays. Such indirect intervention will not affect galaxy evolution in a short timescale but could be substantial over long cosmic time. Our observations prove galactic magnetic fields are already present in the Cosmic Dawn, which supports the picture that galactic magnetic fields could co-evolve with the host galaxy over a considerable cosmic period.
A quickly formed magnetic field could also intervene with the earliest galaxy formation. James Webb Space Telescope (JWST) has recently reported an overabundance of massive galaxies at (Harikane et al. 2023; Finkelstein et al. 2023; Labbé et al. 2023). The possible explanations proposed so far include Pop-III stars (e.g. Ventura et al. 2024), top-heavy initial mass functions (IMFs) (e.g. Yung et al. 2024), feedback-free starbursts (Dekel et al. 2023), early positive AGN feedback (e.g. Silk et al. 2024) or simply the incorrect modelling of their stellar masses (e.g. Wang et al. 2024). The quick formation of proto-galaxies also indicates that various dynamos could start much earlier and quicker, leading to an early formation of G magnetic fields in the ISM. We, therefore, propose that magnetic fields could be another viable factor to impact the formation of early massive galaxies, including the processes such as enhancing the formation of massive stars and modifying the feedback process of the starburst and AGN (e.g. Sharda et al. 2020; Sharda & Menon 2024; Stacy et al. 2022; Sadanari et al. 2024; Begelman & Silk 2023). In the future, surveys of magnetic fields in more distant galaxies could provide better statistical evidence about the importance of magnetic fields in the early Universe.
Meanwhile, the rapid formation of galactic magnetic fields could also play an important role on a larger scales. A statistical study of nearly 200 nearby galaxies has found that magnetic field strength decreases as a function of distance from the galactic disk, but it is enhanced in the bipolar direction, indicating that the CGM is magnetised by the galactic outflow (Heesen et al. 2023). With sensitive low-frequency telescopes, like Low Frequency Array (LOFAR), magnetic fields have been observed up to Mpc around the galaxy cluster (Di Gennaro et al. 2020; Cuciti et al. 2022). The early development of magnetic fields in the ISM supports the hypothesis that the magnetic fields could be transferred from the ISM to the CGM (Bertone et al. 2006; Hanasz et al. 2013). This process could be rather efficient, along with the ubiquitous presence of galactic outflows found in dusty star-forming galaxies (e.g. Spilker et al. 2020), which can provide a natural explanation for large-scale scale-magnetic fields in the CGM of local galaxies. polarised dust emission extends over 3 kpc and shows a per cent variance. We measure a bimodal distribution of the magnetic field position angle, peaking at 28 degrees and 78 degrees. The region with polarisation detection is spatially coincident with the [C ii] emission around a velocity component of km s-1, and it also has a generally small velocity gradient and velocity dispersion. Our observations demonstrate that mapping and resolving kpc-scale polarised dust emission back to the Cosmic Dawn is possible with modern submillimetre interferometric observations.
The observed kpc-scale polarised dust emission indicates the presence of large-scale ordered magnetic fields. We discuss the formation and implications of such large-scale ordered magnetic fields. Combining all the available observations, the general properties of SPT034652, and the field structures in the source plane favour a galaxy merger event as the primary driver of the kpc-scale ordered magnetic fields. With the assumption that the observed ordered magnetic fields are stretched from the existing magnetic fields of the pre-merged galaxies, our observation supports kpc-scale ordered magnetic fields can form in the early massive star-forming galaxy by the time the Universe is only 1 Gyr old. Assuming equipartition with the turbulent kinetic energy, we also compute the upper limit of the turbulent magnetic field strength of G, which could be amplified by the extreme star formation activity. The quick formation of magnetic fields provides new implications for forming massive galaxies in the very early Universe. Such early magnetic fields could mediate the early star formation and play a crucial role in the long-term evolution of galaxies. The early presence of magnetic fields in the interstellar medium could also contribute to the formation of large-scale magnetic fields in the circum-galactic medium.
The future ahead is promising. The current results are biased to a single extreme starburst galaxy, but this approach can be easily applied to many other strongly lensed dusty star-forming galaxies (DSFGs) and even the brightest unlensed targets in the early Universe with a reasonable amount of telescope time. Future surveys on a large sample of DSFGs at different redshifts and star formation rates could be used to test whether the magnetic fields are common in the ISM of young star-forming galaxies and how they evolve across cosmic time. In addition, high-resolution observations of the best available targets will reveal more detailed structures on smaller scales, which will be crucial to understanding the possible mechanisms driving their early development.
Acknowledgements.
We thank the anonymous referee for providing useful comments, which significantly improved this manuscript. J.C. is grateful for the useful help from Andy Biggs and Johan Richard, which has inspired this project. J.C. gratefully acknowledges the support provided by European Research Council, grant number 101055023. E.L.-R. is supported by the NASA/DLR Stratospheric Observatory for Infrared Astronomy (SOFIA) under the 08_0012 Program. SOFIA is jointly operated by the Universities Space Research Association,Inc.(USRA), under NASA contract NNA17BF53C, and the Deutsches SOFIA Institut (DSI) under DLR contract 50OK0901 to the University of Stuttgart. E.L.-R. is supported by the NASA Astrophysics Decadal Survey Precursor Science (ADSPS) Program (NNH22ZDA001N-ADSPS) with ID 22-ADSPS22-0009 and agreement number 80NSSC23K1585.This paper makes use of the following ALMA data: 2021.1.00458.S and 2013.1.01231.S. ALMA is a partnership of ESO (representing its member states), NSF (USA), and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile.
[…]